OpenStudy (anonymous):

second question: DOCTOR d has a total of 3000 invested in 2 stocks,stock a dividends at the rate of 6% , stock b at 8%.Her dividends from both 2 stocks is 220 per year,how much did she invest in both stock a and b

OpenStudy (anonymous):

Letting a = stock a and b = stock b a + b = 3000 .06a + .08b = 220 a = 3000- b .06(3000 - b) + .08b = 220 180 - .06b + .08b = 220 .02b = 40 b = 2000 a = 1000